Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used iRequired: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overhComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComRequired 1 Required 2 Required 3 Required 4 Required 5 Prepare a schedule of cost of goods manufactured for the year. AssumeComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 Ass
Answer #1
1 Predetermined overhead rate=Estimated manufacturing overhead/Estimated direct materials used=121800/87000=$ 1.40 per direct material
2 Actual manufacturing overhead:
$
Indirect labor 125600
Property taxes 8700
Depreciation of equipment 19000
Maintenance 14000
Insurance 9300
Rent, building 34000
Total 210600
Manufacturing overhead applied=Actual direct materials used*Predetermined overhead rate
Actual direct materials used:
$
Beginning raw materials 24000
Add: Purchase of raw materials 134000
158000
Less: Ending raw materials 10000
Actual direct materials used 148000
Manufacturing overhead applied=148000*1.40=$ 207200
MOH applied < Actual MOH. Hence,overhead is said to be under-applied.
Under-applied overhead=Actual MOH-Applied MOH=210600-207200=$ 3400
3 Schedule of cost of goods manufactured
$ $
Direct materials:
Raw materials inventory,beginning 24000
Add: Purchase of raw materials 134000
Total raw materials available 158000
Deduct: Raw materials inventory,ending 10000
Raw materials used in production 148000
Direct labor 83000
Manufacturing overhead applied 207200
Total manufacturing costs 438200
Add: Beginning work in process inventory 47000
485200
Deduct: Ending work in process inventory 39000
Cost of goods manufactured 446200
4 Schedule of cost of goods sold
$
Finished goods inventory,beginning 73000
Add: Cost of goods manufactured 446200
Cost of goods available for sale 519200
Deduct: Finished goods inventory,ending 56000
Unadjusted cost of goods sold 463200
5 Work in process inventory=Direct materials+Direct labor+Manufacturing overhead
Manufacturing overhead applied=Actual direct materials used*Predetermined overhead rate=8800*1.4=$ 12320
39000=8800+Direct labor+12320
Direct labor=39000-8800-12320=$ 17880
$
Direct materials 8800
Direct labor 17880
Manufacturing overhead 12320
Work in process inventory 39000